OpenStudy (anonymous):

solve the given equation log2x=4

OpenStudy (anonymous):

2

OpenStudy (yolo_king):

2 * ? = 4

OpenStudy (anonymous):

-16 c. 8 16 d. 4

OpenStudy (anonymous):

it 2 because 2x2=4

OpenStudy (anonymous):

but if the chooses are -16 8 16 or 4

OpenStudy (campbell_st):

raise both sides are powers of the base, its one of the 1st log laws if \[\log_{a} (x) = b~~~~then~~~~~ a^{\log_{a}(x)} = a^b~~~so~~x = a^b\]

OpenStudy (campbell_st):

so in your question you have a = 2 and b = 4

OpenStudy (campbell_st):

looking at the choices... you can eliminate -16 as you can't take the log of a negative.

OpenStudy (anonymous):

\[\log_{2} x=4\]

OpenStudy (yolo_king):

@campbell_st could you help out here after wards http://openstudy.com/updates/54dcf719e4b0b0ad8853e8cf

OpenStudy (anonymous):

my bad i am new and i did not know that there way the equation that i could do